# Johan Håstad

### v2

@@ 4 The switching lemma
Proof:  We are interested in a $\sigma$ that gives a long path in the decision tree. As in previous papers [10], [11] we explore what has been called “the canonical decision tree” under the partial restriction $\rho$ which we from now on call simply “a restriction” dropping the word “partial”.
@@
For any variable $x_{e}$ we define its influential mini-square(s). This is either a single mini-squares or two mini-squares. We want the property that if we know the values of all $y_{P}$ around the influential mini-square(s) then this uniquely determines the value of $x_{e}$. If $e$ is within a mini-square then this mini-square is its influential mini-square(s). If $e$ is in a brick outside the mini-squares then the influential mini-squares are the closest end point(s) of the live path(s) in this brick. By “closest” we here mean a live mini-square such that if we know the value of all adjacent variables $y_{P}$ then we know the value of the $x_{e}$. We do not consider the value of $x_{e}$ known unless we completely know the situation at its influential mini-square(s). If such a mini-square is alive the extra information needed is provided by something we call a matched pair.
Definition 4.2
A matched pair $(s_{1},s_{2})$ is the information that $s_{1}$ and $s_{2}$ should be matched by changing the type of an augmenting path from 0 to 1 between the two mini-squares.
@@
We use this concept for pairs of neighboring chosen mini-squares and for pairs in the matching $\pi_{2}$. In both cases the identity of the path to change follows from the context. For pairs of chosen mini-squares it is the chosen path and for pairs in $\pi_{2}$ it is the path the was modified when it was first picked.
A matched pair $(s_{1},s_{2})$ fixes, once and for all, the values of all variables $y_{P}$ where $P$ is a path with an end-point $s_{1}$ or $s_{2}$
@@
We now proceed to define the canonical decision tree. The process is guided by $\rho$ and a set $I$ of matched pairs. The support of the set $I$ is the mini-squares appearing in any matched pair and this is initially empty.
@@
Letting $T_{i}$ denote $\varphi(F_{i})$, we go over the branches of $T_{i}$ for increasing values of $i$ and in a fixed order for each tree. The forceable branch is the first branch leading to a one that can be followed given the current $I$ and $\rho$. Let us be formal.
@@
Before stage $j$ we have information set $I^{j}$ in the form of some matched pairs. It contains some pairs from $\pi_{2}$ and some pairs based on answers in the decision tree. In stage $j$ we have forcing information $J_{j}$ that makes us follow the forceable branch and it contains.
@@
For any chosen mini-square mentioned in $J_{j}$ we ask for its partner in the decision tree. This information, jointly with the matched pairs from $\pi_{2}$ in $J_{j}$, forms the $j$th information set, $I_{j}$, and we set $I^{j+1}=I^{j}\cup I_{j}$.
Given $I^{j+1}$ we can determine whether the forceable branch is followed. If it is, we answer 1 in the canonical decision tree and halt the process. Otherwise we go to the next stage and look for the next forceable branch. If there is no more forceable branch we halt with answer 0. Let $T$ be the resulting decision tree. To see that this is a an acceptable choice for $\varphi(F)$, we have a pair of lemmas.
Lemma 4.3
@@
Proof:  Suppose there is a locally consistent branch in $T_{i}\lceil_{\sigma\gamma}$ that leads to a leaf labeled one. The information used to follow this branch can be used as forcing information.
Lemma 4.4
@@
Proof:  In fact for $T_{i}$ used for the construction of the forceable branch we have now reached a leaf that is labeled 1.
We use Razborov’s labeling argument [18] to analyze the probability that we make $2s$ queries in the canonical decision tree. Take any branch of this length and suppose it was constructed during $g$ stages using information sets $J_{j}$. Let $J^{*}=\cup_{j=1}^{g}J_{j}$ and as any query is a result of including an element in $J^{*}$ we know that it contains at least $s$ pairs and for notational convenience we assume that this number is exactly $s$. Note that $J^{*}$ may not be locally consistent when seen as a partial matching on the chosen nodes, but this is not required. We proceed to analyze the probability that the process results in a $J^{*}$ of this size. Let us start with an easy observation.
Lemma 4.5
@@
Proof:  The parts coming from $\pi_{2}$ are clearly disjoint as $\pi_{2}$ is a matching. The pairs of chosen nodes are also disjoint as any mini-square included in $J_{j}$ is included in $I_{j}$ and later $J_{j^{\prime}}$ are disjoint from $I_{j}$.
@@
In the spirit of [10] and [11] we want to find a restriction $\rho^{*}$ with fewer live mini-squares and then show how to reconstruct $\rho$ from $\rho^{*}$ and some external information. First note that it is possible for several tuples $(\tau,U,\pi_{2},B)$ to produce the same $\rho$. For this reason we construct a quadruple $\tau^{*},U^{*},\pi_{2}^{*},B^{*}$ and use this together with external information to reconstruct $\tau,U,\pi_{2},B$. The quadruple $\tau^{*},U^{*},\pi_{2}^{*},B^{*}$ corresponds to a restriction $\rho^{*}$.
@@
Informally we want $\rho^{*}$ to be $\rho$ modified by the information set $J^{*}$. In other words for a matched pair in $J^{*}$ we change the type of one augmenting path between the two mini-squares and now consider the two endpoints to be dead. We proceed to find a quadruple that gives this restriction.
@@
As a first approximation we let $\pi_{2}^{*}$ be $\pi_{2}$ with all matched pairs in $J^{*}$ removed and $U^{*}=U$. We need to process the matching pairs, $(s_{1},s_{2})$ of chosen nodes in $J^{*}$. We first remove these two mini-squares from $U^{*}$. If there is a pair $(s_{1}^{\prime},s_{2}^{\prime})$ in $\pi_{2}^{*}$ where $s_{i}^{\prime}$ is in the same super-square as $s_{i}$ we remove this pair from $\pi_{2}^{*}$ and put $s_{1}^{\prime}$ and $s_{2}^{\prime}$ in to $U^{*}$. If there is no such edge we allow for two “holes” of two empty super-squares in $U^{*}$.
After we have constructed $U^{*}$ and $\pi^{*}_{2}$ we now construct $\tau^{*}$ and $B^{*}$ to complete the quadruple. Let us start by setting $\tau^{*}$ to take the values given by $\rho$.
Let us, again in ±$\pm 1$ sum up the number the neighboring variables $y_{P}$ around any mini-square. Due to the changing of the variables corresponding to $\pi_{1}$ and $\pi_{2}$ the sum for any live mini-square, except the designated survivor is $-2$ while the sum around the designated survivor is 0. The sum around the dead mini-squares is 0.
For pairs from $\pi_{1}$ or $\pi_{2}$ that remains matched either in $\pi_{2}^{*}$ or $\pi_{1}^{*}$ we restore the value of $\tau^{*}$ to take the value of $\tau$ and we let the corresponding bits of $B^{*}$ equal to those of $B$.
For pairs in $J^{*}\cap\pi_{2}$ we change the value of the paths used in its construction from 0 to 1 in $\tau^{*}$. We set the bits $B^{*}$ to the value in $B$.
For the pairs of chosen mini-squares in $J^{*}$ we change the value of the $y_{P}$ corresponding to the chosen path from 00 to $1$, we set the values of $B^{*}$ equal to the bits of $B$.
For mini-squares in $U^{*}$ that either are new (because they came from $\pi_{2}$) or has lost their partner (due to the partner belonging to a pair in $J^{*}$) we proceed as follows. If it has a partner in $\pi^{*}$ pick one possible path $P$ between it and its partner and change the value of $y_{P}$ from 0 to 1 in $\tau^{*}$. Set the bits of $B^{*}$ so that this would have been the path picked by $\pi_{1}$.
We have created an almost complete matching with some exception. If the designated survivor belonged to a pair of $J^{*}$ its sum is now +2.
Any mini-square in $U^{*}$ without a partner in $\pi^{*}_{1}$ has sum -2. This can either be the replacement of the designated survivor in its square or due to a hole in $U^{*}$.
@@
We continue to define a process that, using external information, reconstructs $\tau$, $U$, $\pi_{2}$ and $B$ from $\tau^{*}$, $U^{*}$, $\pi_{2}^{*}$ and $B^{*}$ and the trees $T_{i}$. The “star quadruple” determines the values of of $y_{P}$ in the same way as the non-stared quadruple determined values for $y_{P}$. To be more precise the process is the following.
@@
  • For any pair of nodes in $\pi_{1}$ (as a matching of super-squares) find path connecting this pair as indicated by $B^{*}$. Change the corresponding value from $1$ to 00.
@@
Call this restriction $\rho^{*}$. Let us first the key fact.
Lemma 4.6
@@
Proof:  We claim that any variable $y_{p}$ forced by $\rho$ is also forced by $\rho^{*}$ and to the same value. This follows as any such variable requires all its influential mini-squares to be dead. Any mini-square that this is dead in $\rho$ is also dead in $\rho^{*}$. As the situation around it has not changed it is forced to the same value.
The additional information to follow the first forceable branch is given by $J_{1}$. As this information is included in $\rho^{*}$ the first forceable branch is followed.
To correctly identify the first forceable branch is an important step in reconstruction. Unfortunately there might be earlier paths forced to one by $\rho^{*}$. The reason is that the information that makes us follow such a path might come from several different $J_{i}$ and there is no guarantee that these matchings are locally consistent and hence it might not constitute an acceptable set $J_{1}$.
The optimistic view is that any such path still would help us as it must point to some live mini-squares. We are, however, not able to use this and we need to make sure that the reconstruction is not confused. We introduce the concept of “signature” to enable us to identify the first forceable branch.
Definition 4.7
The signature of a live mini-square determines whether it is chosen and in such a case which direction it is matched in its forceable branch.
The reconstruction is now as follows. We start with $\rho^{*}$ and we reconstruct $I_{i}$ and $J_{i}$ in order. We let $\rho_{j}^{*}$ be the restriction obtained from $\rho$ jointly with $I_{i}$ for $i<j$ and $J_{i}$ for $i\geq j$. Thus $\rho^{*}_{1}$ is simply $\rho^{*}$ which is the starting point. We have a set $E$ of prematurely found chosen mini-squares jointly with their signatures. This is initially the empty set. We proceed as follows.
@@
    Read a bit to determine whether there are more live variables to be found on this branch. In such a case, read a number in $[t]$ to determine its position. If this variable has two influential mini-squares, read another two bits to determine which of these two mini-squares are alive. For any alive influential mini-square, retrieve the corresponding signature(s) from external information and, if chosen, include the mini-squares in $E$. Go to step 2.
@@
Lemma 4.8
If the information from $E$ used on the branch leading to a one is consistent with the information in $I^{j-1}$ then it is the forceable branch.
@@
Proof:  Suppose $v\in E$, which by definition, is a chosen mini-square, and was included in $J_{j^{\prime}}$ for some $j^{\prime}\geq j$. As the current path is forced to one it must have been a potential forceable branch. If it was not the actual forceable branch it must be that at least one of the matched pairs needed to follow this branch was not allowed. Forcing information from $\pi_{2}$ is always allowed and thus the only problem could have been consistency on the chosen mini-squares. If none of the elements used on the current path gives any conflict with $I^{j-1}$, it was allowed and hence it must be the forceable branch.
@@
We need to discuss how to reconstruct $I_{j}$ and $J_{j}$ and start with the latter. For each matched pair in $J_{j}$ we have recovered at least one end-point from the forceable branch. If needed we use external information to recover the other end-point. This costs at most $\Delta$. Whenever we recover a pair of adjacent mini-squares we use external information to recover the advice bits. This only costs 2 bits and thus we below focus on identifying mini-squares and do not mention the reconstruction of advice bits.
@@
For each element in $J_{j}$ we need to discover whether it is chosen (this is one bit of external information) and to which node is is matched in $I_{j}$. If it is not chosen then the information in $I_{j}$ is the same as the one in $J_{j}$. If it is chosen then the partner is either in $J_{j}$ and easy to find or it is live in $\rho^{*}_{j}$. This follows as if it is chosen and belongs to $I_{j}$ but not $J_{j}$ then, by Lemma 4.5 it cannot belong to $J^{*}$. In the case when it is alive we can, at cost $2C\log n$, find its identity. Once we have identified the two partners in a matching, at cost $O(1)$ we can reconstruct which augmenting path was used.
@@
We reconstruct $I_{j}$ and $J_{j}$ and compute $\rho_{j+1}^{*}$ and proceed to the next stage. At the end of this process we have reconstructed $\rho$ and in the process we have also identified all the sets $I_{j}$ and $J_{j}$ and we know which pairs in $J_{j}$ are chosen and which belong to $\pi_{2}$. For the pairs of chosen mini-squares we put them back in to $U$ and any replacements are moved back to $\pi_{2}$. At cost $O(1)$ we can identify the advice bits of any single mini-square involved in a move. This way we recover $U$, the full $\pi_{2}$, and $B$. Once we have these we can read off the original $\tau$.
Let us calculate how much information was used. For each edge in $J^{*}$ at least one end-point is discovered at cost at most $t$ and the other at cost at most $\Delta$. On top of this we have information such as the whether it is chosen and the signature at cost $O(1)$. This gives a total cost of at most $2^{O(s)}(t\Delta)^{s}$. In reconstructing $I_{i}$ we might incur an additional cost of $O(\log n)$ per mini-square for an additional factor of $2^{O(s)}(\log n)^{2s}$ and finally a factor $2^{O(s)}$ to get the advice bits for nodes that move out and into $U$. Thus the total cost for the external information is bounded by $2^{O(s)}(\log n)^{2s}(t\Delta)^{s}$.
@@
We proceed to compare the number of quadruples $\tau,U,\pi_{2},B$ to the number of quadruples $\tau^{*},U^{*},\pi_{2}^{*},B^{*}$, and let us first assume that there are no holes. Both $U$ and $U^{*}$ contain one mini-square from each super-square but there is a difference that $U$ contains the designated survivor while $U^{*}$ might contain a different mini-square from this super-square. Thus the number of possibilities for $U^{*}$ is a factor $\Delta$ larger than the number of possibilities for $U$.

### v3

@@ 4 The switching lemma
Remark 4.2.
A good way to think about the parameters when applying Lemma 4.1 is that $R$ is given by the choice of $n$. We then choose $\Delta$ to be sufficiently large to get a small failure probability. The value of $T$ that controls how much the instance shrinks is then calculated as $T=150\Delta^{2}R$, and this also determines the value of $k$.
Proof 4.3.
@@
We are interested in a $\sigma$ that gives a long path in the decision tree. As in previous papers [10], [12] we explore what has been called “the canonical decision tree” under the partial restriction $\rho$ which we from now on call simply “a restriction” dropping the word “partial”. As $\sigma$ has fewer live variables compared to $\rho$ this is sufficient to establish the lemma.
@@
For any variable $x_{e}$ we define its influential mini-square(s). This is either a single mini-squares or two mini-squares. We want the property that if we know the values of all $y_{P}$ around the influential mini-square(s) then this uniquely determines the value of $x_{e}$. If $e$ is within a mini-square then this mini-square is its influential mini-square(s). If $e$ is in a brick outside the mini-squares then the influential mini-squares are the closest end point(s) of the live path(s) in this brick. By “closest” we here mean a live mini-square such that if we know the value of all adjacent variables $y_{P}$ then we know the value of the $x_{e}$.
Definition 4.4.
The value of $x_{e}$ is forced iff the values of $y_{P}$ is known for all paths connected to its influential mini-square(s).
If we are interested in the value of a variable $x_{e}$ that is not forced, we need to find out more information to change this state of affairs. The main way to get such information is by something we call matched pairs.
Definition 4.5.
If two mini-squares $s_{1}$ and $s_{2}$ give a matched pair then we should change the value of $y_{P}$ from 0 to $1$ where $P$ is a path with end-points $s_{1}$ and $s_{2}$.
Remark 4.6.
@@
We use this concept for pairs of neighboring chosen mini-squares and for pairs in the matching $\pi_{2}$. In both cases the identity of the path to change follows from the context. For pairs of chosen mini-squares it is the chosen path and for pairs in $\pi_{2}$ it is the path the was modified when it was first picked. Note that a matched pair $(s_{1},s_{2})$ fixes, once and for all, the values of all variables $y_{P^{\prime}}$, where $P^{\prime}$ is a path with an end-point $s_{1}$ or $s_{2}$
We now proceed to define the canonical decision tree. Remember that while our starting trees have variables corresponding to edges in the grid, the tree we are creating has variables corresponding to the chosen paths.
@@
The process of creating the canonical decision tree is guided by $\rho$ and a set $I$ of matched pairs. The support of the set $I$ is the mini-squares appearing in any matched pair and this is initially empty.
@@
Letting $T_{i}$ denote $\varphi(F_{i})$, we go over the branches of $T_{i}$ for increasing values of $i$. As the values to some variables are forced by the current information we cannot freely follow any branch. We locate the first (in any fixed order) branch that leads to one and which can be followed by the current information and call it the forceable branch. Let us be formal.
@@
Before stage $j$ we have an information set $I^{j}$ in the form of some matched pairs. It contains some pairs from $\pi_{2}$ and some pairs based on answers in the decision tree. In stage $j$ we have forcing information $J_{j}$ that forces the values of all variables on the forceable branch leading to a one. This set contains.
@@
We now extend the canonical decision tree by, for any chosen mini-square mentioned in $J_{j}$ we ask for its partner in the decision tree. This information, jointly with the matched pairs from $\pi_{2}$ in $J_{j}$, forms the $j$th information set, $I_{j}$, and we set $I^{j+1}=I^{j}\cup I_{j}$.
Given $I^{j+1}$ we can determine whether the forceable branch is followed. If it is, we answer $1$ in the canonical decision tree and halt the process. Otherwise, we go to the next stage and look for the next forceable branch. If there is no more forceable branch we halt with answer 0. Let $T$ be the resulting decision tree. To see that this is an acceptable choice for $\varphi(F)$, we have a pair of lemmas.
Lemma 4.7.
@@
Proof 4.8.
Suppose there is a locally consistent branch in $T_{i}\lceil_{\sigma\gamma}$ that leads to a leaf labeled one. The information used to follow this branch can be used as forcing information.
Lemma 4.9.
@@
Proof 4.10.
In fact for $T_{i}$ used for the construction of the forceable branch we have now reached a leaf that is labeled $1$.
We use Razborov’s labeling argument [18] to analyze the probability that we make $2s$ queries in the canonical decision tree. Slightly oversimplifying, given a $\rho$ that gives a long path in the canonical decision tree, we create a $\rho^{*}$ with fewer live centers such that we can recover $\rho$ from $\rho^{*}$ and some limited size external information.
Take any branch of this length and suppose it was constructed during $g$ stages using information sets $J_{j}$. Let $J^{*}=\cup_{j=1}^{g}J_{j}$ and as any query is a result of including an element in $J^{*}$ we know that it contains at least $s$ pairs and for notational convenience we assume that this number is exactly $s$. As the estimate for the probability of getting a set $J^{*}$ of size $s+d$ is exponentially decreasing in $d$ we can sum over all values of $d$, formally justifying this assumption.
Note that $J^{*}$ may not be locally consistent when seen as a partial matching on the chosen nodes, but this is not required. We proceed to analyze the probability that the process results in a $J^{*}$ of size $s$. Let us start with an easy observation.
Lemma 4.11.
@@
The parts coming from $\pi_{2}$ are clearly disjoint as $\pi_{2}$ is a matching. The pairs of chosen nodes are also disjoint as any mini-square included in $J_{j}$ is included in $I_{j}$ and later $J_{j^{\prime}}$ are disjoint from $I_{j}$.
@@
In the spirit of [10] and [12] we want to find a restriction $\rho^{*}$ with fewer live mini-squares and then show how to reconstruct $\rho$ from $\rho^{*}$ and some external information. First note that it is possible for several tuples $(\tau,U,\pi_{2},B)$ to produce the same $\rho$ and hence to make counting unambiguous we count such 4-tuples and not partial restrictions. In the same vein we do not reconstruct it from $\rho^{*}$ but rather from a quadruple $(\tau^{*},U^{*},\pi_{2}^{*},B^{*})$ (and some external information). The quadruple $(\tau^{*},U^{*},\pi_{2}^{*},B^{*})$ does produce a restriction $\rho^{*}$ but that is not what we count. As the value of $\pi_{1}$ is fixed, $\pi_{1}^{U}$ is determined by $U$ and hence we do not need to include $\pi_{1}$ in the information.
@@
Informally we want $\rho^{*}$ to be $\rho$ modified by the information set $J^{*}$. In other words for a matched pair in $J^{*}$ we change the type of one augmenting path between the two mini-squares and now consider the two endpoints to be dead. We proceed to find a quadruple that gives this restriction. The main crux is to construct $\pi_{2}^{*}$ and $U^{*}$ and let us start with this.
@@
We initialize $\pi_{2}^{*}$ to $\pi_{2}$ with all matched pairs in $J^{*}$ removed and $U^{*}=U$. We process the matching pairs, $(s_{1},s_{2})$ of chosen nodes in $J^{*}$ one by one in the order as they appear in $J_{j}$. If there is a pair $(s_{1}^{\prime},s_{2}^{\prime})$ in $\pi_{2}^{*}$ where $s_{i}^{\prime}$ is in the same super-square as $s_{i}$ we remove this pair from $\pi_{2}^{*}$ and replace $s_{i}$ by $s_{i}^{\prime}$ in $U^{*}$ for $i=1$ and $2$. If there is no such edge we allow for two “holes” of two empty super-squares in $U^{*}$ as we still remove $s_{1}$ and $s_{2}$.
Next we construct $\tau^{*}$. There is not any real choice how to do this as we want $\rho^{*}$ to look like $\rho$ with the additional information given by $J^{*}$ but let us go over the details. For any square already dead in $\rho$ nothing changes and all information stays the same. For the live nodes there are a number of cases. One important property to keep in mind is that for an alive but not chosen node the pairings are the same in $\pi_{2}$, $J^{*}$ and their $I_{j}$ sets. When it comes to the chosen nodes this is not the case and they can be matched to different nodes in $\pi_{1}^{U}$, $J^{*}$ and their $I_{j}$ sets. Let us consider a mini-square $s$ in super-square $S$. We have a few alternatives.
  1. 1.
    The mini-square $s$ is contained in a matched pair from $\pi_{2}\cap J^{*}$.
  2. 2.
    The mini-square $s$ is contained in a matched pair from $\pi_{2}$ and moved to $U^{*}$.
  3. 3.
    The mini-square $s$ is contained in a matched pair from $\pi_{2}$ and is not moved to $U^{*}$ and does not belong to $J^{*}$.
  4. 4.
    The mini-square $s$ contained in $U$ and neither it nor its partner in $\pi_{1}^{U}$ takes part in the above process.
  5. 5.
    The mini-square $s$ is contained in $U$ and not in a matched pair in $J^{*}$, but where its partner in $\pi_{1}^{U}$ is contained in a pair in $J^{*}$.
  6. 6.
    The mini-square $s$ is contained in $U$ and is matched in $J^{*}$.
The values of $\tau^{*}$ and $\tau$ are the same at most points and we use “changing the value” to indicate the value is different in the two mappings.
Case 1.
We have that $s$ is now dead in $\rho^{*}$. We change one value by restoring the value when $s$ was introduced to $\pi_{2}$.
Case 2. First note that as the pairs in $J^{*}$ are disjoint, $s$ remains in $U^{*}$ for the duration of the process. If its partner from $\pi_{2}$ is in the super-squares which is the partner of $S$ in the matching $\pi_{1}$ then $\tau^{*}$ equals $\tau$ around $s$. If it is not then we change the value at one or two position. If $s^{\prime\prime}$ is the partner in $\pi_{2}$ restore the value of the variable of the path between $s$ and $s^{\prime\prime}$ when $(s,s^{\prime\prime})$ was put into $\pi_{2}$. If $s$ is in the super-square of the designated survivor do nothing more and otherwise consider the super-square, $S^{\prime}$, that is the partner of $S$ under $\pi_{1}$. If there is a node $s^{\prime}\in S^{\prime}$ in $U^{*}$ change the value of one path variable between $s$ and $s^{\prime}$ from 1 to 0. We choose one variable that is possible through our selection function.
Case 3.
This is simple as $\tau^{*}$ equals $\tau$ around $s$.
Case 4.
Also in this case $\tau^{*}$ equals $\tau$ around $s$.
Case 5.
Suppose $s^{\prime\prime}$ is the partner in $\pi_{1}^{U}$ that disappeared. We restore the value of the chosen path between $s^{\prime\prime}$ and $s$ to 1. Let $S^{\prime}$ be the partner of $S$ under $\pi_{1}$. If there is a node $s^{\prime}\in S^{\prime}$ in $U^{*}$ change the value of one path variable between $s$ and $s^{\prime}$ from 1 to 0. We choose one variable that is possible through our selection function.
Case 6.
In the case $s$ is now dead in $\rho^{*}$. We restore the value on the path used when constructing $\pi_{1}^{U}$.
This completes the description of $\tau^{*}$. Note that it might not be exactly an almost complete matching as, due to the possibility of holes in $U^{*}$ and the possibility that the element in the super-square of the designated survivor is no more the designated survivor. This causes an unbalance of one at some mini-squares, but this deviation is completely determined by $U^{*}$.
The value of $B^{*}$ plays little role. We simply set it to equal $B$. We might have to change it for pairs of mini-squares participating in the above process, but these are few and this can be handled by the external information.
@@
We now proceed to define a process that, using external information, reconstructs the tuple $(\tau,U,\pi_{2},B)$ from $(\tau^{*},U^{*},\pi_{2}^{*},B^{*})$ and the trees $T_{i}$. The “star quadruple” determines the values of $y_{P}$ in the same way as the non-stared quadruple determined values for $y_{P}$. To be more precise the process is the following.
@@
  • For any pair of mini-squares in $\pi_{1}^{U^{*}}$ find one path connecting this pair as indicated by $B^{*}$. Change the corresponding value from $1$ to 0.
@@
Call this restriction $\rho^{*}$. Let us establish a simple lemma for warm-up just to give an indication of how the proof proceeds.
Lemma 4.13.
@@
Proof 4.14.
We claim that any variable $y_{p}$ forced by $\rho$ is also forced by $\rho^{*}$ and to the same value. This follows as any such variable requires all its influential mini-squares to be dead. Any mini-square that this is dead in $\rho$ is also dead in $\rho^{*}$. As the situation around it has not changed it is forced to the same value.
The additional information needed to follow the first forceable branch is given by the matching pairs in $J_{1}$. As this information is included in $\rho^{*}$ the first forceable branch is followed.
To correctly identify the first forceable branch is an important step in the reconstruction. Unfortunately there might be earlier branches forced to one by $\rho^{*}$. The reason is that the information that makes us follow such a branch might come from several different $J_{i}$ and there is no guarantee that these matchings are locally consistent and hence it might not constitute an acceptable set $J_{1}$.
The optimistic view is that any such branch still would help us as it must point to some live mini-squares. We are, however, not able to use this and we need to make sure that the reconstruction is not confused. We introduce the concept of “signature” to enable us to correctly identify the first forceable branch.
Definition 4.15.
The signature of a live mini-square determines whether it is chosen and in such a case which direction it is matched in its forceable branch. It is given by three bits.
Let us turn to the reconstruction process. During this process we specify the needed external information and in each situation we give the number of bits needed. We sum up the total size of the external information at the end.
We start with $\rho^{*}$ and we reconstruct $I_{j}$ and $J_{j}$ in order. We let $\rho_{j}^{*}$ be the restriction obtained from $\rho$ jointly with $I_{i}$ for $i<j$ and $J_{i}$ for $i\geq j$. In particular, $\rho^{*}_{1}$ is simply $\rho^{*}$ which is the starting point. We have a set $E$ of prematurely found chosen mini-squares jointly with their signatures. This is initially the empty set. We proceed as follows.
@@
    Read a bit to determine whether there are more live variables to be found on the current branch. In such a case, read a number in $[t]$ to determine its position. If this variable has two influential mini-squares, read another two bits to determine which of these two mini-squares are alive. For any alive influential mini-square, retrieve the corresponding signature(s) from external information and, if chosen, include the mini-squares in $E$. Go to step 2.
@@
Lemma 4.16.
If the mini-squares of $E$ together with their signatures on a branch leading to a one is consistent with the information in $I^{j-1}$, then it is a forceable branch. In particular the first such branch after the j1j-1th forceable branch is the $j$th forceable branch.
Proof 4.17.
@@
Suppose $v\in E$, which by definition, is a chosen mini-square, and was included in $J_{j^{\prime}}$ for some $j^{\prime}\geq j$. As the current path is forced to one it must have been a potential forceable branch. If it was not the actual forceable branch it must be that at least one of the matched pairs needed to follow this branch is not allowed. Forcing information from $\pi_{2}$ is always allowed and thus the only possible problem is the consistency on the chosen mini-squares. If the signatures of the mini-squares used on the current branch do not give any conflict with $I^{j-1}$, it was allowed and hence it must be a forceable branch.
@@
Whenever we recover a pair of adjacent mini-squares we use external information to recover the advice bits. This only costs 2 bits and thus we below focus on identifying mini-squares and do not mention the reconstruction of advice bits.
We need to discuss how to reconstruct $I_{j}$ and $J_{j}$ and start with the latter. For each matched pair in $J_{j}$ we have recovered at least one end-point from the forceable branch. If needed we use external information to recover the other end-point. This costs at most $1+\log\Delta$ bits. This recovers $J_{j}$ and we need to extend it to $I_{j}$.
@@
For each element in $J_{j}$ we need to discover whether it is chosen (this is one bit of external information) and to which node it is matched in $I_{j}$. If it is not chosen then the information in $I_{j}$ is the same as the one in $J_{j}$ (and also in $\pi_{2}$). If it is chosen then the partner is either in $J_{j}$ and easy to find or it is live in $\rho^{*}_{j}$. This follows as if it is chosen and belongs to $I_{j}$ but not $J_{j}$ then, by Lemma 4.11 it cannot belong to $J^{*}$. In the case when it is alive we can, at cost $O(1)+\log\log n$ bits, find its identity. Once we have identified the two partners in a matching, at cost $O(1)$ we can reconstruct which augmenting path was used. This reconstructs all of $I_{j}$.
@@
We reconstruct $I_{j}$ and $J_{j}$ and compute $\rho_{j+1}^{*}$ and proceed to the next stage. At the end of this process we have reconstructed $\rho$ and in the process we have also identified all the sets $I_{j}$ and $J_{j}$ and we know which pairs in $J_{j}$ are chosen and which belong to $\pi_{2}$. For the pairs of chosen mini-squares, we put them back in to $U$ and any replacements are moved back to $\pi_{2}$. At cost $O(1)$ we can identify the advice bits of any single mini-square involved in a move. This way we recover $U$, the full $\pi_{2}$, and $B$. Once we have these we can read off the original $\tau$.
Let us calculate how much information was used. We have the following contributions.
  • For many mini-square processed we need to read the signature. As we only process $O(s)$ mini-squares this is $O(s)$ bits.
  • For the chosen nodes, once we have established their identities we might need $O(1)$ bits to determine in which direction they are matched. Note that this can be different in $J_{j}$ and $I_{j}$. This costs at most $O(s)$ bits.
  • For potentially forceable branches we read one bit to determine if there is any additional live mini-square to be found through this branch. As there are only at most $s$ forceable branches the answer can be “no” at most $s$ times. As we only discover at most $O(s)$ mini-squares to put in to $E$, the answer can only by “yes” at most $O(s)$ times. Thus the total number of such bits read is $O(s)$.
  • For each edge in $J^{*}$ at least one end-point is discovered at cost at most $1+\log t$ bits and the other at cost at most $1+\log\Delta$ bits.
  • For each mini-square in $I_{j}$ that need be discovered since it was not a member of $J_{j}$ we might need $\log\log n+O(1)$ bits.
Summing up the information used we first have a cost $O(s)$ in several (but constant number) places. The main cost is given by the construction of the pairs in $J^{*}$ and this is $s(\log t+\log\Delta)$. For each mini-square in $J^{*}$ we might reconstruct its partner in $I_{j}$ at an additional total cost at most $2s\log\log n$.
@@
We proceed to compare the number of quadruples $(\tau,U,\pi_{2},B)$ to the number of quadruples $(\tau^{*},U^{*},\pi_{2}^{*},B^{*})$, and let us first assume that there are no holes. Both $U$ and $U^{*}$ contain one mini-square from each super-square but there is a difference that $U$ contains the designated survivor while $U^{*}$ might contain a different mini-square from this super-square. Thus the number of possibilities for $U^{*}$ is a factor $\Delta$ larger than the number of possibilities for $U$.